### v3

# Development of the cadmium zinc TElluride Radiation Imager (TERI)

Abstract
The cadmium zinc TElluride Radiation Imager, or TERI, is an instrument to space qualify large-volume $4\times 4\times 1.5\ \mathrm{cm}^{3}$ pixelated CdZnTe (CZT) detector technology. The CZT’s anode is composed of a $22\times 2$ 22 array of pixels while the cathode is planar. TERI will contain four of those crystals with each pixel having an energy range of $40\ \mathrm{keV}$ up to $3\ \mathrm{MeV}$ with a resolution of $1.3\%$ full-width-at-half maximum at $662\ \mathrm{keV}$ all while operating in room temperature. As the detectors are 3D position sensitive, TERI can Compton image events. TERI is fitted with a coded-aperture mask which permits imaging low energy photons in the photoelectric regime. TERI’s primary mission is to space-qualify large-volume CZT and measure its degradation due to radiation damage in a space environment. Its secondary mission includes detecting and localizing astrophysical gamma-ray transients. TERI is manifested on DoD’s STP-H10 mission for launch to the International Space Station in early 2025.
keywords:
CdZnTe, Gamma-ray imaging, gamma-ray spectroscopy, gamma-ray astronomy, coded-aperture imaging
1 INTRODUCTION
[…]
DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.
To address the gap in astronomical observations of MeV gamma rays, several missions are being developed and conceptualized utilizing various technologies [1]. Observations in the MeV range will aid in understanding the formation, evolution, and physics of astrophysical jets, which will identify the physical processes in the extreme conditions around compact objects and the study of the life cycle of matter in the nearby universe. In light of this, we are developing the cadmium zinc TElluride Radiation Imager (TERI) to increase its technological readiness for potential future use of space-based CdZnTe (CZT) sensors.
CZT is by far not new to operating in the space environment. NASA’s Swift-BAT Observatory [2] has operated planar CZT since 2004 (and is still operational as of February 2024). Coplanar CZT flew on the NASA Dawn spacecraft for an interplanetary mission. Moreover, pixelated CZT with a volume of $4\times 4\times 0.5\ \mathrm{cm}^{3}$ flew on several Indian Space Research Organization instruments including AstroSat and RT-2 [3, 4]. This is the largest crystal we have found that has flown yet.
We are therefore fielding TERI with four $4\times 4\times 1.5\ \mathrm{cm}^{3}$ crystals, the largest CZT crystal developed to date [5]. A larger single-volume crystal is desired to result in a higher detection efficiency. A larger single volume reduces the guard-ring-to-volume ratio thereby increasing the detection efficiency [6]. In addition, larger crystals could also have a positive effect on overall spectroscopic performance. Pixels near the edges generally perform poorly due to the poorer electric field. A larger crystal again reduces the perimeter-to-volume ratio. The improvements originating from larger crystals will enhance the overall efficiency of the imager-spectrometer instrument.
COSI is a small explorer class telescope in development to survey the soft MeV sky (0.2-5 MeV) [7]. Its technology is based on double-sided double strip detector technology which uses high purity germanium (HPGe) as its detection medium. Due to the nature of HPGe, it needs to be cooled to cryogenic temperatures. This adds complexities to the system, which adds cost and engineering requirements. Room temperature semiconductors, such as CZT, relax the temperature requirements while maintaining good resolution compared to scintillator based instruments. The TERI detectors therefore offer an alternative to HPGe.
TERI is currently planned for a launch in early 2025 to the International Space Station (ISS) on the Department of Defense (DoD) Space Test Program (STP) STP-H10 mission. It is slated to be hosted on SpaceX Commercial Resupply Services CRS-32. This manuscript serves as its instrument paper. Sec. 2 will overview the entire instrument. Sec. 3 described its in-lab spectroscopic performance, Sec. 4 describes the coded aperture, while Sec. 5 presents its Compton imaging capabilities.
2 TERI Overview
The cadmium zinc TElluride Radiation Imager, or TERI, has a total volume of $35\mathrm{(L)}\times 35\mathrm{(W)}\times 31\mathrm{(H)}\ \mathrm{cm}^{3}$. It weighs $\sim 40\ \mathrm{kg}$ and has a power draw of $\sim 30\ \mathrm{W}$. Fig. 1 shows an exploded view of TERI. There are two major components, the electronics housing and the mask assembly. The electronics housing contains the 4 CZT detectors, a power distribution unit, a single board computer (Raspberry Pi) to act as the master flight computer, and an Ethernet switch. The upper half of the instrument consists of the coded mask assembly, which is discussed in Sec. 4. Fig. 2(a) shows a cross-sectional view of TERI while Fig. 2(b) shows the fully assembled instrument.
DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.Refer to captionDISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited.DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Coded Mask DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Shadow Shield DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Closeout Panels DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Mask Assembly DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Electronics Box DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Ethernet Switch DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Single Board Computer DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Power Distribution Board DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Support Ribs DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. 4x H3D Detector ‘Bricks’ DISTRIBUTION STATEMENT A. Approved for public release: distribution is unlimited. Spacecraft Interface
Figure 1: Exploded view of the TERI (computer aided design model). The mechanical interface to the Columbus CEPA pallet is not depicted.
Table 1: Table with TERI’s instrument performance. The unshielded field of view (FOV) refers to the opening of the shadow shield. The Compton imaging resolution is calculated with a $\mathrm{Cs}$-137 source placed in the center of a single TERI detector. The reported value reflects the FWHM of the reconstructed point spread function (PSF). The Maximum Likelihood Expectation Maximization (MLEM) value is calculated under similar conditions using 25 MLEM iterations. See Section 5 for additional information on TERI’s Compton imaging capabilities.
Metric Value
Energy Range $40\ \mathrm{keV}-3\ \mathrm{MeV}$ per pixel
[…]
Unshielded FOV $0.69\ \mathrm{sr}$
Compton Imaging Resolution $45^{\circ}\ \mathrm{FWHM\ (SBP)}$ $<12^{\circ}\ \mathrm{FWHM\ (MLEM-25)}$
Coded Aperture Imaging Resolution $<2^{\circ}\ \mathrm{FWHM}$
Volume $35\mathrm{(L)}\times 35\mathrm{(W)}\times 31\mathrm{(H)}\ \mathrm{cm}^{3}$
Mass $\sim 40\ \mathrm{kg}$
Power $\sim 30\ \mathrm{W}$
(a) Cross-Sectional View
(b) TERI Picture
Figure 2: Figures of TERI. (a) Shows a cross-sectional view with the Zenith’s vector roughly pointing to the right. The ISS forward direction is out of the page. (b) Picture of the fully assembled TERI.
2.1 Mission Programmatics
Launch is provided by the Space Test Program (STP), under the United States Space Force, with mission support for at least one year. TERI is one of a few instruments integrated on a Columbus External Payload Adapter (CEPA) pallet designated STP-H10. The pallet’s avionics supplies 28V to the instrument along with a GPS-synchronized pulse per second (PPS) signal. Communication between the pallet and TERI is performed via an asynchronous RS-422. The mission’s payload is slated to occupy Starboard Overhead X (SOX) pointing spot on the Columbus external payload facility. TERI will be pointed $15^{\circ}$ away from the ISS overhead vector towards the ISS starboard to prevent field-of-view obstruction from other ISS structures and visiting vehicles. TERI’s orientation is fixed relative to the ISS and will not alter during flight.
Figure 3: Picture of the International Space Station with SOX and the Columbus module identified. Image is adapted from NASA.gov [8].
2.2 TERI’s CZT Detectors
Each CZT crystal has a volume of $4\times 4\times 1.5\ \mathrm{cm}^{3}$. The anode is pixelated with a $22\times 2$ 22 array and a pixel pitch of $1.77\ \mathrm{mm}$, while the cathode remains planar. This electrode configuration allows for the 3D position estimation of each interaction. We voxelized the interactions using the 2D location of the triggered pixel’s center. The 3rd dimension, or the depth of interaction, is estimated using one of two ways. If a single interaction in the crystal has occurred, we use the signals from the cathode and anode and take their ratio (cathode-to-anode ratio). We will use the drift time to estimate the depth if more than one interaction has occurred in the crystal. The techniques for the depth of interaction estimation are extensively explored in He [9], Kaye [10], and Zhang [11]. We then apply a voxel by voxel energy correction [12].
Fig. 4(a) shows one of the flight CZT crystals. The rest of the detector unit, which we will refer to as a module is manufactured by H3D, Inc., and is a customized version of the T400 model. Due to the nature of the detector readout, coincidence cannot be accomplished between multiple detectors.
The modules are placed in a pinwheel design to keep a compact form factor while maintaining the crystals square to each other, as visible in Fig. 4(b). The center-to-center distance between each crystal is ${}9.9\ \mathrm{cm}$.
(a) $4\times 4\times 1.5\ \mathrm{cm}^{3}$ CdZnTe
(b) Electronics housing
Figure 4: Figures showing the detectors in TERI. (a) Shows the $4\times 4\times 1.5\ \mathrm{cm}^{3}$ CdZnTe. (b) Shows a CAD of the top view of the electronics box showing the 4 detector modules in a pinwheel design. The dimensions of the electronics box is $35\mathrm{(L)}\times 35\mathrm{(W)}\times 16.2\mathrm{(H)}\ \mathrm{cm}^{3}$.
Table 2: Table with the detector characteristics.
Metric Value
CZT Volume $4\times 4\times 1.5\ \mathrm{cm}^{3}$
Pixel Configuration $22\times 2$ 22 array
Pixel Pitch $1.77\ \mathrm{mm}$
2.3 Data and Power
Fig. 5 shows the basic electrical schematic of TERI. The 28V line is provided by the pallet’s avionics tower. It is then filtered, and split into a 28V DC-DC converter and a 5V DC-DC converter. The 28V provides power to the detector modules. The 5V line provides power to a Raspberry Pi (the flight computer) and an Ethernet switch. The Raspberry Pi communicates with each of the detectors via the Ethernet switch. Each detector provides the 3D position of each interaction along with the reconstructed energy and time. It monitors voltages, currents, and temperature probes located within TERI. The housekeeping data along with histogramed and list-mode data from the detectors is telemetered to the ground with the Consultative Committee for Space Data Systems (CCSDS) protocol.
Figure 5: Electrical schematic of the TERI system. The pallet avionics tower provides 28V power.
3 Spectroscopic Performance
The general performance of the $4\times 4\times 1.5\ \mathrm{cm}^{3}$ CZT crystals is extensively reported in Zhu et alli [5]. A 3D voxel-by-voxel calibration is applied for energy reconstruction that is also temperature dependent [13]. The measured full width at half maximum (FWHM) at $662\ \mathrm{keV}$ is $1.3\%$ using all events and $1.2\%$ using single-pixel events. Fig 6a plots the $662\ \mathrm{keV}$ full energy peak using either single pixel or all pixel events. The resolution, as expected, is better for single-pixel events. Fig. 6b plots the FWHM as a function of energy. Note that this plot uses all pixel event types.
Figure 6: (a) Cs-137 spectra demonstrating the difference in resolution between single pixel and all events. (b) FWHM as a function of energy using all events.
Fig. 7 and 8 plot the spectral response of TERI to different check sources: Am-241, Ba-133, Cs-137, Na-22, and Co-60 (with gamma-ray emission statistics available in [14]). The energy range for each pixel is $40\ \mathrm{keV}$ to $3\ \mathrm{MeV}$. This implies that multi-pixel events can result in the detection of gammas with incident energy higher than $3\ \mathrm{MeV}$.
Figure 7: Energy spectra of Am-241 and Ba-133 taken by TERI. All pixel events are plotted.
Figure 8: Energy spectra of Co-60, Cs-137, and Na-22 taken by TERI. All pixel events are plotted.
4 Code Aperture Assembly
TERI is fitted with a rank 54 pseudorandom-coded aperture mask. We chose a pseudorandom mask, rather than utilize a (M)URA type pattern, as TERI’s detectors are sparsely placed. The random noise properties therefore present the appropriate choice rather than a structured mask [15]. To design the pattern, we create a very large sample of random masks. We then simulated the response by convolving the detector plane’s response with the projected pattern and then producing the associated image. We then calculate the mean square error (MSE) of the reconstructed image. This was repeated for source locations uniformly sampled across the field of view. The mask with the smallest overall MSE was chosen.
Fig. 9(a) shows the chosen mask pattern. The mask is made out of 80 mils ($\sim 2\ \mathrm{mm}$) tantalum square tiles. Each tile is $5.31\ \mathrm{mm}$ wide, which is 3 times the pitch of the CZT’s pixel ($1.77\ \mathrm{mm}$) to reduce collimation effects. Fig. 9(b) shows the assembled mask where each tile is epoxied onto an aluminum panel by hand. The mask-to-detector distance is 6.317 inches with an opening angle of 48.9 deg. The resulting coded field of view is 0.69 steradians.
(a) TERI’s Coded Mask Pattern
(b) TERI Mask
Figure 9: Coded Mask of TERI
4.1 Coded Aperture Imaging
This work uses cross-correlation to reconstruct the images, although more advanced filtering or iterative reconstruction algorithms could be applied. The reconstruction formulation is as follows:
$\hat{\textbf{X}}=\textbf{D}\otimes\textbf{Y},$ (1)
with $\hat{\textbf{X}}$ representing the estimated image, D is the decoding matrix, and Y is the recorded shadowgraph with the measured counts in each pixel representing its elements, $\otimes$ represents the cross-correlation operator. Since TERI utilizes a random mask, the mask is its own decoding function [16]. Fig. 10a displays the recorded pattern by TERI’s detector when a Co-57 source is placed in its central axis where a portion of the mask pattern is visible. Using Eq. 1, we can reconstruct the image shown in Fig. 9(b). In the image, a point source is visible in the center. We note that there are visible structured artifacts in the background. Those are most likely due to the poor packing fraction of the detector plane and the resulting sparse recording of the projected pattern.
Figure 10: (a) is the recorded pattern by the TERI detectors (Y in Eq. 1) when a Co-57 source is placed in the central axis. (b) is the reconstructed image ($\hat{\textbf{X}}$ in Eq. 1).
Fig. 11 plots the image presented in Fig. 10 with a Mollweide projection. This projection displays the field of view of the instrument at any given instance.
Figure 11: Mollweide projection of a Co-57 source presented in Fig. 10b. It is of a Co-57 source placed in the central axis.
Fig. 12 shows reconstructed images using TERI where a different number of detectors are turned off. Fig. 12a plots the image reconstructed with 3 detectors off and progressively turn on different detectors all the way through Fig. 12d where no detectors are turned off. Even with only one detector turned on, TERI can reconstruct a clear image of the source, demonstrating the power of utilizing a random coded mask. As more and more detectors are turned on, the signal-to-noise ratio improves. Note that the background artifacts are changing with the different number of operational detectors.
Figure 12: Reconstructed images of a Co-57 source utilizing a different number of detectors. (a) plots the image when 3 detectors are off, (b) 2 are off, (c) 1 is off, while (d) has all detectors operational.
5 Compton Imaging
Compton imaging is a technique to reconstruct the gamma ray’s source position and distribution when the gamma ray’s energy is large enough to where Compton scattering is the dominant mode of interaction [17]. Compton imaging was demonstrated using many different families of semiconductors [18, 19, 20] and scintillators [21].
Briefly, Compton imaging is accomplished if the incident gamma rays undergo Compton scattering followed by a photoelectric absorption. The incident direction of the gamma ray can be determined with the scattering angle between the incident and scattered photons using the Compton scatter equation. The possible locations where the gamma rays can originate are reconstructed as a ring in the image space. As more photons are collected, more rings overlap and eventually converge to the source position and distribution. Compton ring and imaging reconstruction is fully explored in Kierans et alli [22]. As the TERI detectors are 3D position sensitive, the position of interaction can be determined along with the deposited energy.
At the date of composition of this work, the data pipeline for the full TERI instrument is not yet complete and is still under development. However, we present the gamma-ray imaging performance of one of the four detectors on TERI. We use reconstruction techniques as outlined in Xu [23]. This work makes use of two types of reconstruction algorithms. The first is simple backprojection (SBP) which simply projects the Compton rings onto a spherical image space. Since this is known as a biased estimation, we also present maximum likelihood expectation maximization (MLEM) which is an iterative image reconstruction algorithm [24]. Our MLEM reconstruction techniques follow those outlined in Wilderman et alli [25]. We do not implement any advanced iterative stopping criteria and stop the iterative process based on qualitative criteria. The number of interactions per event used in this section range from 2 to 5 [26]. In addition, we apply far-field approximations.
Figure 13(a) features a SBP reconstruction of a single Cs-137 source placed 36 inches (0.91 meters) directly above the normal of the detector, or $(\mathrm{\theta_{azimuth},\phi_{polar}})=(90^{\circ},90^{\circ})$. The image uses 15k photopeak events and results in a point spread function (PSF) of roughly $45^{\circ}$ FWHM. Figure 13(b) applies MLEM for 25 iterations which results in an FWHM of $<12^{\circ}$.
We then complexify the source scene by creating a multi-location dataset. Since we do not have multiple sources of the same intensity, we use the same source and place it in multiple locations. We then collect the data for the same runtime at each location. We place the same source in four different locations in a ‘T’ configuration: the center, $20^{\circ}$ to the left, right, and below the center. Using coordinates, this represents $(90^{\circ},90^{\circ}),(70^{\circ},90^{\circ}),(110^{\circ},90^{\circ}),(90^{\circ},70^{\circ})$. Figure 13(c) plots the SBP of the 4-source scene using 54k events. The PSF of the hotspot does show some extension relative to the single source scene. We then apply 100 MLEM iterations in Figure 13(d) which clearly unveils the 4-source ‘T’ configuration. As the sources are in the near field and we do not apply any near field correction or advance system response, we observe a relative weakening of the offset sources relative to the center source. Application of an appropriate system response will correct for such effects. This briefly presents TERI’s Compton imaging capability for resolving complex sources.
Figure 13: Reconstructed Compton images in equirectangular projections of a Cs-137 source using only 1 TERI detector. The top row presents a single source located at $(90^{\circ},90^{\circ})$ reconstructed using (a) simple backprojection and (b) 25 MLEM iterations. The bottom row shows a complex source field where the same Cs-137 source is placed in 4 different location (‘T’ configuration) and the datasets concatenated to mimic one single observation.
6 Conclusion
We developed the cadmium zinc TElluride Radiation Imager (TERI), an instrument that utilizes four large-volume $4\times 4\times 1.5\ \mathrm{cm}^{3}$ pixelated CdZnTe (CZT). The main objective of TERI is to qualify and characterize the technology for operation in a space environment. Each pixel has an energy range of $40\ \mathrm{keV}$ up to $3\ \mathrm{MeV}$ with a resolution of $1.3\%$ full-width-at-half maximum at $662\ \mathrm{keV}$. TERI is fitted with a coded aperture which allows for low-energy imaging. In this work, we demonstrated the technologies’ coded-aperture imaging and Compton imaging capabilities. TERI is slated for launch to the International Space Station in early 2025 on DoD’s STP-H10 mission.
7 Code and Data Availability
The authors do not have permission to share data.
Acknowledgements.
We thank the previous researchers whose work on CZT made TERI possible. The authors are particularly grateful to the Department of Defense (DoD) Space Test Program (STP) for providing space launch services and mission operations support for TERI. Daniel Shy appreciates Jacob Levinrad for assisting with mask assembly. Special thanks to Scott Budzien for providing the NRL Payload Operations Control Center. The CdZnTe crystals were provided by the Defense Threat Reduction Agency. Daniel Shy is supported by the U.S. Naval Research Laboratory’s Jerome and Isabella Karle Fellowship. This work is supported by the Office of Naval Research 6.1 funds.
Appendix A History of CZT in space based and balloon-borne experiments
Table 3 attempts to plot a comprehensive list of CZT instruments developed specifically for the space and near-space environment. It starts with mostly balloon instruments using thin planar crystals that are on the order of a few millimeters. Over time, the active area of the crystals increased to $4\times 4\ \mathrm{cm}^{2}$ and evolved from planar to pixelated electrode configurations. Nevertheless, the deployed crystals remained under a centimeter of thickness [^1] We note that in certain applications, like X-ray astrophysics, larger crystals may not be desired., which is not ideal for gamma-ray astronomy due to the small interaction probability.
Table 3: Collection of published balloon and space-based instruments that utilize CZT or CdTe. If two names appear in a single entry that is separated with a colon, the first name corresponds to the name of the instruments while the second name is that of the entire probe. The listed instruments are orbital unless stated otherwise. The ‘s’ shorthands is used to designate slated launch dates. Concepts are limited to known active projects. This list attempts to be comprehensive, however, we acknowledge that we might have missed a few.
Name Crystal Size Material Electrode Config. Year Launched Notes
PoRTIA $2.54\times 2.54\times 0.19\ \mathrm{cm}^{3}$ $1.2\times 1.2\times 0.5\ \mathrm{cm}^{3}$ CZT planar 1995 balloon [27]
CfA/CalTech $2.54\times 2.54\times 0.19\ \mathrm{cm}^{3}$ CZT planar 1995 balloon [28]
HEXIS $1.2\times 1.2\times 0.3\ \mathrm{cm}^{3}$ CZT cross strip/planar 1997 balloon [29]
[…]
ISGRI: Integral $0.4\times 0.4\times 0.2\ \mathrm{cm}^{3}$ CdTe planar 2002  [31]
BAT: Swift $0.4\times 0.4\times 0.2\ \mathrm{cm}^{3}$ CZT planar 2005  [2]
HEFT $1.2\times 2.4\times 0.2\ \mathrm{cm}^{3}$ CZT pixelated 2005 balloon [32]
GRaND: Dawn $1\times 1\times 0.7\ \mathrm{cm}^{3}$ CZT coplanar 2007  [33, 34]
HEX: Chandrayaan-1 $4\times 4\times 0.2\ \mathrm{cm}^{3}$ CZT planar 2008  [35]
AAUSat-2 $1\times 1\times 0.4\ \mathrm{cm}^{3}$ CZT pixelated 2008  [36]
ProtoEXIST $1.95\times 1.95\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2009- balloon [37]
RT-2: Koronas-Foton $3.96\times 3.95\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2009 [4]
NuSTAR $2\times 2\times 0.2\ \mathrm{cm}^{3}$ CZT pixelated 2012  [38]
X(L)-Calibur $2\times 2\times 0.2\ \mathrm{cm}^{3}$ $2\times 2\times 0.5\ \mathrm{cm}^{3}$ $2\times 2\times 0.8\ \mathrm{cm}^{3}$ CZT pixelated 2014- balloon [39]
CZTI: AstroSat $3.9\times 3.9\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2015  [3]
BeEagleSat $2\times 2\times 0.5\ \mathrm{cm}^{3}$ CZT cross strip 2017  [40]
EPEx $2\times 2\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2018 balloon [41]
ASIM $2\times 2\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2018  [42]
STIX: Solar Orbiter $1\times 1\times 0.1\ \mathrm{cm}^{3}$ CdTe variable pixelation 2020  [43]
OrionEagle $2\times 2\times 1.5\ \mathrm{cm}^{3}$ CZT pixelated 2021 balloon [44]
BADG3R $2\times 2\times 0.6\ \mathrm{cm}^{3}$ CZT cross strip 2022 balloon [45]
VZLUSAT-2 $1.4\times 1.4\times 0.2\ \mathrm{cm}^{3}$ CdTe pixelated 2022  [46]
Sharjah-Sat 1 $2.54\times 2.54\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated 2023  [47, 48]
ComPair $0.6\times 0.6\times 2\ \mathrm{cm}^{3}$ CZT virtual Frisch grid 2023 balloon [49, 50]
ECLAIRs: SVOM $0.4\times 0.4\times 0.1\ \mathrm{cm}^{3}$ CdTe Schottky-type 2024  [51]
MASS-Cube $2\times 2\times 1.5\ \mathrm{cm}^{3}$ CZT pixelated s2024  [52]
AXIS: AEPEX $2\times 2\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated s2024  [53, 54]
TERI $4\times 4\times 1.5\ \mathrm{cm}^{3}$ CZT pixelated s2025 this work
COSI $8\times 8\times 1.5\ \mathrm{cm}^{3}$ HPGe cross strip s2027  [7]
Daksha $3.9\times 3.9\times 0.5\ \mathrm{cm}^{3}$ CZT pixelated concept  [55]
ASTENA $0.5\times 2\times 2\ \mathrm{cm}^{3}$ CZT strip-drift concept  [56, 57]
MeVCube $2\times 2\times 1.5\ \mathrm{cm}^{3}$ CZT pixelated concept  [58]
GECCO $0.8\times 0.8\times 3.2\ \mathrm{cm}$ CZT virtual Frisch grid concept  [59]
HSP $2\times 2\times 0.3\ \mathrm{cm}^{3}$ CZT pixelated concept  [60]